xx Fortran Interfaces for Flesh Functions

Fortran Interfaces for Flesh Functions

Erik Schnetter <schnetter@aei.mpg.de>

Date

Abstract

This thorn provides Fortran interfaces for the flesh functions.

1 Introduction

2 Physical System 3 Numerical Implementation 4 Using This Thorn 4.1 Obtaining This Thorn 4.2 Basic Usage 4.3 Special Behaviour 4.4 Interaction With Other Thorns 4.5 Examples 4.6 Support and Feedback 5 History 5.1 Thorn Source Code 5.2 Thorn Documentation 5.3 Acknowledgements References 6 Parameters

11one 10Scope: privatex01  INT

11Description: This parameter has the value one

11Range10  x01  Default: 1

111:1 101

11

7 Interfaces

General Implements: fortran

8 Schedule This section lists all the variables which are assigned storage by thorn CactusBase/Fortran. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage NONE

Scheduled Functions

CCTKxPARAMCHECK x checkfortranparameters x test whether fortran parameters work correctly x

11x11Language:11fortran
11 x 11Type: 11function